# revtex4-2Repair the float

### v1

@@ I Introduction
Majorana zero modes (MZMs) emerge as zero-energy edge states in topological superconductors, a topological phase that has been predicted to occur in superconductor-semiconductor systems at strong Zeeman fields [1, 2, 3]. Since its conception, this idea has motivated an impressive amount of theoretical and experimental works searching for Majorana signatures [4, 5, 6, 7] but it is still unclear whether MZMs have been observed or not [8]. The main issue was shown to be the proliferation of trivial zero-energy states (TZESs) well below the expected topological phase transition [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], with signatures similar to those due to MZMs [21, 22, 23, 24, 25, 26]. This in turn has motivated intense efforts aiming at distinguishing between MZMs and TZESs [27, 28, 29, 19, 30, 31] but the problem is far from being settle [8].
In this work we show that the Majorana polarization, which characterizes the expectation value of the particle-hole operator, can be used as a real space topological indicator that distinguishes between MZMs and TZESs. The Majorana polarization was used before to characterize MZMs but in the absence of TZESs [32, 33, 34, 35]. We extend the previous definition and introduce the Majorana polarization as quantity that captures the the correlation between MZMs emerging at the edges. To show the usefulness of the Majorana polarization, we first consider a finite superconductor with spin-orbit coupling under the presence of a Zeeman field and demonstrate that it detects the topological phase transition and the emergence of MZMs. In this case, we find that the Majorana polarization exactly captures the behavior of the Pfaffian, which then enables us to interpret the Majorana polarization as a real space topological indicator. Notably, we discover that the Majorana polarization remains a good topological invariant even under strong scalar disorder, where the visibility of the gap closing is degraded and an avalanche of disorder-induced TZESs emerge.
@@ II Models and method
$\begin{split}\mathcal{H}=&\sum_{j=1,\sigma\sigma^{\prime}}^{M}d_{j\sigma}^{\dagger}\left[\varepsilon_{\mathrm{}}(j)\delta_{\sigma\sigma^{\prime}}+B\sigma_{\sigma\sigma^{\prime}}^{x}+i\Delta(j)\sigma_{\sigma\sigma^{\prime}}^{y}\right]d_{j\sigma^{\prime}}\\ &+\sum_{j=1,\sigma\sigma^{\prime}}^{M-1}d_{j\sigma}^{\dagger}\left[-t_{\mathrm{}}\delta_{\sigma\sigma^{\prime}}+\alpha_{\mathrm{}}\sigma_{\sigma\sigma^{\prime}}^{y}\right]d_{j+1,\sigma^{\prime}}+\text{ H.c. },\end{split}$ (1)
where $d_{j\sigma}$ and $d^{\dagger}_{j\sigma}$, respectively, destroy and create a fermionic state with spin $\sigma$ at site $x$ in the superconducting system and runs over the entire system of length $L=Ma$, being $M$ the total number of sites and $a$ the lattice spacing. Also, $\varepsilon_{\mathrm{}}(j)=2t_{\mathrm{}}-\mu(j)+V_{\mathrm{}}(j)$ is the onsite energy, with $t$ being the nearest neighbor hopping amplitude, while $\mu(j)$ is the spatial dependent chemical potential and $V(j)$ a spatial dependent potential characterizing scalar disorder, which will be specified later. Moreover, in Eq. (1) $\alpha_{\mathrm{}}$ is the SOC strength, $B$ the Zeeman field along $x$, and $\Delta(j)$ the space dependent superconducting pair potential. Thus, the finite superconductor in Fig. 1(a) of length $L$ is described by an homogeneous pair potential $\Delta(j)=\Delta$ and finite chemical potential $\mu(j)=\mu$; the potential $V(j)$, however, still depends on space, see below. In this single superconducting nanowire, a topological phase transition occurs at $B=B_{\rm c}$ where $B_{\rm c}=\sqrt{\Delta^{2}+\mu^{2}}$. Then, MZMs emerge when $B>B_{\rm c}$.
@@
We argue that the issue with the Pfaffian in inhomogenous systems can be resolved by using the Majorana polarization, which, as we show in this work, turns out to be a more powerful topological indicator because it allows to distinguish between MZMs and TZESs. For this reason, we start by defining the Majorana polarization at each site $x$ as [33]
$P_{j}\equiv\frac{\langle\Psi_{j}|\mathcal{C}|\Psi_{j}\rangle}{\langle\Psi_{j}|\Psi_{j}\rangle}=\frac{\sum_{\sigma=\uparrow,\downarrow}2u_{j\sigma}^{(0)}\big{[}v_{j\sigma}^{(0)}\big{]}^{*}}{\sum_{\sigma=\uparrow,\downarrow}\big{|}u_{j\sigma}^{(0)}\big{|}^{2}+\big{|}v_{j\sigma}^{(0)}\big{|}^{2}}\,,$ (3)
where $\mathcal{C}=\tau^{x}\otimes\sigma^{0}$ is the particle-hole operator, with $\tau^{i}$ and $\sigma^{i}$ the $i$-th Pauli matrix in Nambu and spin spaces, respectively, while $u_{j\sigma}^{(0)}$ and $v_{j\sigma}^{(0)}$ represent the electron and hole components with spin $\sigma$ of the wavefunction $\Psi_{j}$ at site $j$ of the lowest energy state. At this point, we anticipate that in a finite topological superconductor, the Majorana polarization $P_{j}$ of the left and right halves of the system develop opposite signs. We show this effect by taking a two-site Kitaev chain at the sweet spot where MZMs appear, see Appendix A. Motivated by this fact, we take Eq. (3) and define the Majorana polarizations associated to the left and right half of the system, which we refer to as left (L) and right (R) Majorana polarizations and obtain them as,
@@
$\mathcal{P}=P_{\rm L}\cdot P_{\rm R}\,,$ (5)
which we refer to as nonlocal Majorana polarization because it simultaneously takes into account the behaviour of $P_{\rm L/R}$. Since $P_{L/R}$ are normalized, the maximum value they can take is $P_{L/R}=\pm 1$, thus placing the nonlocal Majorana polarization within the range $\mathcal{P}\in[-1,1]$: A value of $\mathcal{P}=-1$ implies that the MZMs are true zero-modes and do not overlap, whereas a slight deviation from $\mathcal{P}=-1$ corresponds that there is cross-talk between the MZMs or that there is a finite energy splitting of the MZMs energy. Any other value of $\mathcal{P}$ signals a non Majorana state, namely, a topologically trivial state. With all these characteristics, it is expected that $\mathcal{P}$ not only senses the topological transition but also measures the correlation between MZMs emerging at the edges of the system and, therefore, able to detect the Majorana spatial nonlocality. Before going further, it is worth noting that, while the local Majorana polarization (3) is similar to the definition adopted in Refs. [40, 33], the nonlocal polarization introduced here [Eq. 5] goes beyond previous considerations because it accounts for the behaviour on the two halves of the system  [^1] It is worth mentioning that the Majorana polarization has been also explored in the many-body context [52, 53, 54, 55, 56], with a definition that differs but aims to characterize the same properties of MZMs. Here we will explore the nonlocal Majorana polarization $\mathcal{P}$, which will be obtained for an open chain in real space.
@@ III.1 Superconducting nanowire without disorder
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a)(b)(c)(d)
@@
The vanishing (finite) value of the Majorana polarization in the trivial (topological) phase seen in Fig. 2(c) can be understood by inspecting the real space dependence of the onsite Majorana polarization contributions, denoted by $P_{j}$ in Eq. (3). In Fig. 3 we plot $P_{j}$ as a function of space at Zeeman fields in the trivial and topological phase, indicated by orange and magneta bars in Fig. 2(c). In the trivial phase $B<B_{\rm c}$, the onsite Majorana polarization $P_{j}$ oscillates in space acquiring amplitudes that are opposite in sign, which, when summing them up to obtain the left and right Majorana polarizations $P_{\rm L/R}$ in Eq. (4), give vanishing contribution for the Majorana polarization $\mathcal{P}$ in Eq. (5). In contrast, in the topological phase $B>B_{\rm c}$, the onsite Majorana polarization $P_{j}$ develops large values at the edges and exhibits an oscillatory decay with amplitudes having the same sign up to the middle of the system $L/2$: $P_{j}<0$ for the left half, while $P_{j}>0$ for the right half. Thus, when summing up $P_{j}$ to find the left and right Majorana polarizations $P_{\rm L/R}$ in Eq. (4), $P_{\rm L/R}$ do not vanish, in contrast to what occurs in the trivial phase. As a result, the finite value of both $P_{\rm L}$ and $P_{\rm R}$ prevent the overall vanishing value of $\mathcal{P}=P_{\rm L}P_{\rm R}$ in Eq. (5), which then explains why $\mathcal{P}$ is finite in the topological phase. When normalizing it to the density-of states, we find $P_{\rm L}=-1$ and $P_{\rm R}=1$, which gives $\mathcal{P}=-1$. Before closing this part, we point out that the behaviour of the Majorana polarization follows the corresponding profile of the spin projection, as we indeed observe in Fig. 3(c,d) for the spin density along $z$, obtained as $S^{z}=\Psi^{\dagger}_{j}\sigma_{z}\Psi_{j}$, where $\Psi_{j}$ is the wavefunction associated to the lowest energy state. Of particular interest is that the onsite Majorana polarization follows the same sign of the spin projection, revealing that the negative sign of $\mathcal{P}$ has natural physical interpretation. In sum, the Majorana polarization is not only a good real space topological indicator that signals the topological phase transition but it can also be used to identify the formation of MZMs and their spatial overlap.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a)(b)(c)(d)(e)(f)(g)(h)(i)
@@ III.2 Superconducting nanowire with disorder
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a)(b)(c)(d)
@@ IV SN junctions with trivial zero-energy states
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a)(b)(c)(d)(e)(f)Refer to captionRefer to captionRefer to captionRefer to caption
@@ Appendix A Majorana polarization in a two-site Kitaev chain
where we have used the subscript L/R because $\Psi_{E^{0}_{+/-}}$ are located in the left/right site. At this point we remark that the left and right Majorana state exhibit a Majorana polarization with opposite sign. Then, by defining a nonlocal Majorana polarization as $p=p_{\rm L}\cdot p_{\rm R}$, it is natural to interpret that $p=-1$ provides information about the existence of both Majorana-like states perfectly localized in the left and right sites of the two-site Kitaev model given by Eq. (7) in the sweet spot. Although there is no exponential protection, no topological number, the total Majorana polarization is able to identify the formation of the two MZMs. For obvious reasons, these MZMs have been coined poor man’s Majorana modes [50].
Away from the sweet spot, when $\varepsilon_{\rm L/R}=\varepsilon$ and $\Delta\neq t$, the Majorarana polarizations become $p_{\rm L}=-p_{\rm R}={\Delta}/{\sqrt{\Delta^{2}+\varepsilon^{2}}}$, which then leads to $p=p_{\rm L}\cdot p_{\rm R}\approx-1+2({\varepsilon}/{\Delta})^{2}$ for $\varepsilon\ll\Delta$. Hence, as $\varepsilon$ moves away from zero acquiring finite values, the nonlocal Majorana polarization becomes $p<-1$ signaling that MZMs hybridize. For $\varepsilon\gg\Delta$, the Majorana polarization reads $p\approx-(\Delta/\varepsilon)^{2}$, which practically becomes a very small number, again signalling a strong Majorana hybridization. In the general case, $\varepsilon_{\rm L}\neq\varepsilon_{\rm R}$, MZMs are challenging to realize and the nonlocal Majorana polarization $p$ takes a convoluted form.
@@ Appendix B Comparison between the Pfaffian and Majorana polarization in disordered systems
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a)(b)(c)(d)(e)(f)(g)(h)

### v2

# Identifying trivial and Majorana zero-energy modes using the Majorana polarization

@@ I Introduction
Majorana zero modes (MZMs) emerge as zero-energy edge states in topological superconductors, a topological phase that has been predicted to occur in superconductor-semiconductor systems at strong Zeeman fields [1, 2, 3, 4]. Since its conception, this idea has motivated an impressive amount of theoretical and experimental works searching for Majorana signatures [5, 6, 7, 8, 9, 10] but it is still unclear whether MZMs have been observed or not [11]. The main issue was shown to be the proliferation of trivial zero-energy states (TZESs) well below the expected topological phase transition [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], with signatures similar to those due to MZMs [25, 26, 27, 28, 29, 30]. This in turn has motivated intense efforts aiming at distinguishing between MZMs and TZESs [31, 32, 33, 22, 34, 35] but the problem is far from being settle [11].
In this work we show that the Majorana polarization, which characterizes the expectation value of the particle-hole operator, can be used as a real space topological indicator that distinguishes between MZMs and TZESs in superconducting systems [Fig. 1]. The Majorana polarization was used before to characterize MZMs but in the absence of TZESs [36, 37, 38, 39, 40]. The Majorana polarization was initially defined as an absolute value function and involved only over one half of the system. As such, this initial definition of the Majorana polarization is unable to distinguish between MZMs and TZESs and cannot capture the nonlocal correlations between MZMs (Majorana nonlocality) which involves having both ends of the system. We extend the previous definition and introduce the Majorana polarization as quantity that captures the correlation between MZMs emerging at the edges, and eliminate the use of absolute value. To show the usefulness of the Majorana polarization, we first consider a finite superconductor with spin-orbit coupling under the presence of a Zeeman field and demonstrate that it detects the topological phase transition and the emergence of MZMs. In this case, we find that the Majorana polarization exactly captures the behavior of the Pfaffian, which then enables us to interpret the Majorana polarization as a real space topological indicator. Notably, we discover that the Majorana polarization remains a good topological invariant even under strong scalar disorder, where the visibility of the gap closing is degraded and an avalanche of disorder-induced TZESs emerge.
@@ II Models and method
$\begin{split}\mathcal{H}=&\sum_{j=1,\sigma\sigma^{\prime}}^{M}d_{j\sigma}^{\dagger}\left[\varepsilon_{\mathrm{}}(j)\delta_{\sigma\sigma^{\prime}}+B\sigma_{\sigma\sigma^{\prime}}^{x}+i\Delta(j)\sigma_{\sigma\sigma^{\prime}}^{y}\right]d_{j\sigma^{\prime}}\\ &+\sum_{j=1,\sigma\sigma^{\prime}}^{M-1}d_{j\sigma}^{\dagger}\left[-t_{\mathrm{}}\delta_{\sigma\sigma^{\prime}}+\alpha_{\mathrm{}}\sigma_{\sigma\sigma^{\prime}}^{y}\right]d_{j+1,\sigma^{\prime}}+\text{ H.c. },\\ \end{split}$ (1)
where $\varepsilon(j)=2t-\mu(j)+V(j)$ is the site-dependent onsite energy with $t$ being the nearest neighbor hopping amplitude, while $\mu(j)$ is the spatial dependent chemical potential and $V(j)$ a spatial dependent potential characterizing scalar disorder, which will be specified later. Here $d_{j\sigma}$ and $d^{\dagger}_{j\sigma}$, respectively, destroy and create a fermionic state with spin $\sigma$ at site $j$ in the superconducting system and runs over the entire system of length $L=Ma$, being $M$ the total number of sites and $a$ the lattice spacing. Moreover, in Eq. (1) $\alpha_{\mathrm{}}$ is the SOC strength, $B$ the Zeeman field along $x$, and $\Delta(j)$ the space dependent superconducting pair potential. Thus, the finite superconductor in Fig. 1(a) of length $L$ is described by an homogeneous pair potential $\Delta(j)=\Delta$ and finite chemical potential $\mu(j)=\mu$; the potential $V(j)$, however, still depends on space, see below. In this single superconducting nanowire, a topological phase transition occurs at $B=B_{\rm c}$ where $B_{\rm c}=\sqrt{\Delta^{2}+\mu^{2}}$. Then, MZMs emerge when $B>B_{\rm c}$, where the system becomes a $p$-wave superconductor and effectively describes the physics of the Kitaev chain [41].
@@
We argue that the issue with the Pfaffian in inhomogenous systems can be resolved by using the Majorana polarization, which, as we show in this work, turns out to be a more powerful topological indicator because it allows to distinguish between MZMs and TZESs. For this reason, we start by defining the Majorana polarization at each site $j$ as [37]
$P_{j}\equiv\frac{\langle\Psi_{j}|\mathcal{C}|\Psi_{j}\rangle}{\langle\Psi_{j}|\Psi_{j}\rangle}=\frac{\sum_{\sigma=\uparrow,\downarrow}2u_{j\sigma}^{(0)}\big{[}v_{j\sigma}^{(0)}\big{]}}{\sum_{\sigma=\uparrow,\downarrow}\big{|}u_{j\sigma}^{(0)}\big{|}^{2}+\big{|}v_{j\sigma}^{(0)}\big{|}^{2}}\,,$ (3)
where $\mathcal{C}=\tau^{x}\otimes\sigma^{0}\mathcal{K}$ is the particle-hole operator, with $\mathcal{K}$ the conjugation operator, $\tau^{i}$ and $\sigma^{i}$ the $i$-th Pauli matrix in Nambu and spin spaces, respectively, while $u_{j\sigma}^{(0)}$ and $v_{j\sigma}^{(0)}$ represent the electron and hole components with spin $\sigma$ of the wavefunction $\Psi_{j}$ at site $j$ of the lowest energy state. At this point, we anticipate that in a finite topological superconductor, the Majorana polarization $P_{j}$ of the left and right halves of the system develop opposite signs. We show this effect by taking a two-site Kitaev chain at the sweet spot where MZMs appear, see Appendix A. Motivated by this fact, we take Eq. (3) and define the Majorana polarizations associated to the left and right half of the system, which we refer to as left (L) and right (R) Majorana polarizations and obtain them as,
@@
$\mathcal{P}=P_{\rm L}^{*}\cdot P_{\rm R}=P_{\rm L}\cdot P_{\rm R}^{*}\,,$ (5)
which we refer to as nonlocal Majorana polarization because it simultaneously takes into account the behaviour of $P_{\rm L/R}$. Note that, while $P_{\rm L/R}$ is in general complex, $\mathcal{P}$ is real. Since $P_{\rm L/R}$ are normalized, the maximum absolute value they can take is unity, thus placing the nonlocal Majorana polarization within the range $\mathcal{P}\in[-1,1]$: A value of $\mathcal{P}=-1$ implies that the MZMs are true zero-modes and do not overlap, whereas a slight deviation from $\mathcal{P}=-1$ corresponds that there is cross-talk between the MZMs or that there is a finite energy splitting of the MZMs energy. Any other value of $\mathcal{P}$ signals a non Majorana state, namely, a topologically trivial state. With all these characteristics, it is expected that $\mathcal{P}$ not only senses the topological transition but also measures the correlation between MZMs emerging at the edges of the system and, therefore, able to detect the Majorana spatial nonlocality. Following a similar spirit, the nonlocal Majorana polarization given by Eq. (5) can be easily generalized to higher dimensions, in the same way as the local Majorana polarization [38, 37]. Before going further, it is worth noting that, while the local Majorana polarization (3) is similar to the definition adopted in Refs. [38, 37], the nonlocal polarization introduced here [Eq. 5] goes beyond previous considerations because it accounts for the behaviour on the two halves of the system  [^1] It is worth mentioning that the Majorana polarization has been also explored in the many-body context [60, 61, 55, 62, 63, 64], with a definition that differs but aims to characterize the same properties of MZMs. Similarly, the Majorana polarization has been used to characterize higher order topological superconductors [65]. We note that our definition in Eqs. (4) and (5) has now been adopted to distinguishing between MBSs and other zero-energy modes in quasicrystals [66]. Here we will explore the nonlocal Majorana polarization $\mathcal{P}$, which will be obtained for an open chain in real space.
@@ III.1 Superconducting nanowire without disorder
The vanishing (finite) value of the Majorana polarization in the trivial (topological) phase seen in Fig. 2(c) can be understood by inspecting the real space dependence of the onsite Majorana polarization contributions, denoted by $P_{j}$ in Eq. (3). In Fig. 3 we plot $P_{j}$ as a function of space at Zeeman fields in the trivial and topological phase, indicated by orange and magneta bars in Fig. 2(c). In the trivial phase $B<B_{\rm c}$, the onsite Majorana polarization $P_{j}$ oscillates in space acquiring amplitudes that are opposite in sign, which, when summing them up to obtain the left and right Majorana polarizations $P_{\rm L/R}$ in Eq. (4), give vanishing contribution for the Majorana polarization $\mathcal{P}$ in Eq. (5). In contrast, in the topological phase $B>B_{\rm c}$, the onsite Majorana polarization $P_{j}$ develops large values at the edges and exhibits an oscillatory decay with amplitudes having the same sign up to the middle of the system $L/2$: $P_{j}<0$ for the left half, while $P_{j}>0$ for the right half. Thus, when summing up $P_{j}$ to find the left and right Majorana polarizations $P_{\rm L/R}$ in Eq. (4), $P_{\rm L/R}$ do not vanish, in contrast to what occurs in the trivial phase. As a result, the finite value of both $P_{\rm L}$ and $P_{\rm R}$ prevent the overall vanishing value of $\mathcal{P}=P_{\rm L}P_{\rm R}^{*}=P_{\rm L}^{*}P_{\rm R}$ in Eq. (5), which then explains why $\mathcal{P}$ is finite in the topological phase. When normalizing it to the density-of states, we find $P_{\rm L}=-1$ and $P_{\rm R}=1$, which gives $\mathcal{P}=-1$. Before closing this part, we point out that the behaviour of the Majorana polarization follows the corresponding profile of the spin projection, as we indeed observe in Fig. 3(c,d) for the spin density along $z$, obtained as $S^{z}=\Psi^{\dagger}_{j}\sigma_{z}\Psi_{j}$, where $\Psi_{j}$ is the wavefunction associated to the lowest energy state. Of particular interest is that the onsite Majorana polarization follows the same sign of the spin projection, revealing that the negative sign of $\mathcal{P}$ has natural physical interpretation. In sum, the Majorana polarization is not only a good real space topological indicator that signals the topological phase transition but it can also be used to identify the formation of MZMs and their spatial overlap.
@@ VI Concluding remarks
The nonlocal Majorana polarization investigated here also leaves open questions. For instance, the fact that the Majorana polarization is able to identify the topological phase transition and the formation of Majorana zero modes suggests that it captures properties of the emergent topological superconducting state. As such, there seems to be an intriguing relation between Majorana polarization and emergent superconductivity, which might also require to generalize the definition adopted here to the many-body realm [55] and therefore deserves further investigation. Another interesting problem is about the measurement of the Majorana polarization. To access the Majorana polarization, it is necessary to obtain the coherence factors $u$ and $v$, see Eq. (3). One possible approach to measure such coherence factors is through Andreev spectroscopy [56], which can be complemented by measurements of the local density of states measurements [57]. Although this task is nontrivial, detecting the Majorana polarization would help characterizing the nature of zero-energy states in Majorana devices.
@@ Appendix A Majorana polarization in a two-site Kitaev chain
where we have used the subscript L/R because $\Psi_{E^{0}_{+/-}}$ are located in the left/right site. At this point we remark that the left and right Majorana state exhibit a Majorana polarization with opposite sign. Then, by defining a nonlocal Majorana polarization as $p=p_{\rm L}^{*}\cdot p_{\rm R}=p_{\rm L}\cdot p_{\rm R}^{*}$, it is natural to interpret that $p=-1$ provides information about the existence of both Majorana-like states perfectly localized in the left and right sites of the two-site Kitaev model given by Eq. (7) in the sweet spot. Although there is no exponential protection, no topological number, the total Majorana polarization is able to identify the formation of the two MZMs. For obvious reasons, these MZMs have been coined poor man’s Majorana modes [58].
Away from the sweet spot, when $\varepsilon_{\rm L/R}=\varepsilon$ and $\Delta\neq t$, the Majorarana polarizations become $p_{\rm L}=-p_{\rm R}={\Delta}/{\sqrt{\Delta^{2}+\varepsilon^{2}}}$, which then leads to $p=p_{\rm L}\cdot p_{\rm R}\approx-1+2({\varepsilon}/{\Delta})^{2}$ for $\varepsilon\ll\Delta$. Hence, as $\varepsilon$ moves away from zero acquiring finite values, the nonlocal Majorana polarization becomes $p>-1$ signaling that MZMs hybridize. For $\varepsilon\gg\Delta$, the Majorana polarization reads $p\approx-(\Delta/\varepsilon)^{2}$, which practically becomes a very small number, again signalling a strong Majorana hybridization. In the general case, $\varepsilon_{\rm L}\neq\varepsilon_{\rm R}$, MZMs are challenging to realize and the nonlocal Majorana polarization $p$ takes a convoluted form.